1. At ABC University, a group of students was selected and asked how much of their weekly allowance they spent in buying mobile phone load. The following is the list of amounts spent: Php 120, 110, 100, 200, 10, 90, 100,

100. Calculate the mean, the range, and the standard deviation.



1
Expert's answer
2022-01-10T23:53:47-0500

Sample mean:

"\bar x=\frac{1}{n}\Sigma_{i=1}^nx_i=\frac{830}{8}=103.75."

Sample standard deviation:

"s=\sqrt{\frac{1}{n-1}\Sigma_{i=1}^n(x_i-\bar x)^2}=\sqrt{\frac{18587.5}{7}}\approx51.53."

Range:

"r=x_{max}-x_{min}=200-10=190."